OpenStudy (anonymous):

Find f(x) and g(x) so the function can be expressed as y = f(g(x)). y = Two divided by x squared. + 3

OpenStudy (anonymous):

Is is this : \[\frac{2}{x^2+3}\] or this : \[\frac{2}{x^2} +3\]

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

f(x)=\[2divx +3\] g(x)=\[x^{2}\] hence y=the given answer

OpenStudy (anonymous):

no, its \[\frac{ 2 }{ x^2}+3\]

OpenStudy (anonymous):

but f and g are the above functions and y is a composite of them

OpenStudy (anonymous):

no im saying that for the equation, y= \[\frac{ 2 }{ x^2}+3\]

OpenStudy (anonymous):

yes ure right but the question is findin f and g right?

OpenStudy (anonymous):

yes, they want the function expressed as y=f(g(x)).